OpenStudy (peterg2001):

Choose the answer based on the most efficient method as presented in the lesson. If the first step in the solution of the equation -9 + x = 5x - 7 is "subtract x," then what should the next step be? subtract 5x add 7 add 9

OpenStudy (peterg2001):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

Same thing... -9 + x = 5x - 7 -9=4x-7 Which makes more sense: 0=4x+2 or -2=4x.